# Robust $M$-Estimation of Scatter Matrices via Precision Structure Shrinkage

### v1

@@ 2.4 Penalized $M$-estimation
which is the negative log-likelihood divided by $n$. The function $\rho^{*}(s)\coloneqq-\log g(s)$ ($s\geq 0$) is defined for $s\geq 0$.
@@
and seek $V\in\mathcal{S}_{++}^{p}$ that minimizes objective function (2). In this case, the minimizer $V$ of (2) is an $M$-Estimator [20, 23, 36]. For example, TME corresponds to the choice $\rho(s)=p\log s$ ($s>0$) in (2) [36]. In this paper, no assumptions are imposed on the function $\rho(\cdot)$ itself. Instead, assumptions are placed on the function corresponding to its first derivative, as stated in Condition 2.1.
@@ 2.5 Characterization of the estimator
A solution satisfying the estimating equation (4) is linearly equivariant (Remark 2), whereas a solution satisfying (5) is not, and this difference affects the robustness of the estimator [6]. In (3), when the Kullback–Leibler type penalty
@@
for $\alpha\in(0,\infty)$ is considered, the estimating equation is obtained as
@@
The $M$-estimator of $V$ is often defined through estimating equations such as (4), (5), and (6), and computed using methods based on fixed-point algorithms for the corresponding estimating equations. The estimator characterized as solutions to the estimating equations (4), (5), and (6) do not necessarily possess scale equivariance. See Remark 2 for the definition of scale equivariance.
@@
    The weigt function $w:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ is nonincreasing.
@@
    For $\psi(s)\coloneqq sw(s)$ ($s\geq 0$), it holds that $\kappa\coloneqq\sup_{s}\{\psi(s)\}<\infty$.
@@
    The function $\psi:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ is strictly increasing in $\{s\in\mathbb{R}_{\geq 0};\psi(s)<\kappa\}$.
@@
    For any (p1)(p-1)-dimensional linear subspace $H$, it holds that $P(H)<1-p/\kappa$.
@@
for $c>0$, the $t$-distribution type weight functions $w(s)=(p+\nu)/(s+\nu)$ ($s\geq 0$) for $\nu\in\mathbb{N}$, and Tyler-type weight function $w(s)=p/s$ ($s\geq 0$), which leads to distribution-free properties within the class of elliptical distributions.
2.6 Additive finite-sample breakdown point
The breakdown point is one of the measures of an estimator’s robustness against contamination by outliers in the data [14, 21, 39, 40].
@@ 2.6 Additive finite-sample breakdown point
where $\mathcal{E}_{m}(\mathrm{X})=\{\mathrm{Z}=\mathrm{X}\cup\mathrm{Y};\,\hat{\theta}(\mathrm{Z})\textnormal{ exists in }\Theta\,(\mathrm{X}:\text{fixed})\}$.
@@
$\varepsilon^{*}(X;\,\hat{V}(\cdot),\mathrm{X})\coloneqq\min\left\{\varepsilon_{m};\,b\left(\varepsilon_{m};\,\hat{V}(\cdot),\mathrm{X}\right)=\infty\right\}.$
We consider a discrepancy function $B(V_{1},V_{2})$ that satisfy
$B(V_{1},V_{2})\to\infty\quad\text{as }\lambda_{p}(V_{2})\to 0\ \text{or}\ \lambda_{1}(V_{2})\to\infty.$
Examples of such discrepancy functions include $\mathrm{Tr}[V_{1}V_{2}^{-1}+V_{1}^{-1}V_{2}]$ and $\lVert\log(V_{1}^{-1/2}V_{2}V_{1}^{-1/2})\rVert_{\mathrm{F}}$ [39, 40]. Here, for $X\in\mathcal{S}_{++}^{p}$, we define
@@
where $Q\,\mathrm{diag}\left(\lambda_{1}(X),\dots,\lambda_{m}(X)\right)\,Q^{\top}$ is the eigenvalue decomposition of $X$ with an orthogonal matrix $Q$.
@@ 3.1 Limitations of existing methods
We examine the performance of TME and LNSMI through numerical experiments under two types of outlier configurations. We observe that when outliers form clusters the performance of TME and LNSMI deteriorates significantly for a sample size of $N=100$ and an outlier ratio of $3\%$, when the data dimension exceeds approximately $p\approx 35$. See Section 5 for the detailed settings and Figure 3-(b, d, f) for the result. Moreover, this phenomenon is found to be closely related to the breakdown point of TME. Based on the results of the numerical experiments, we conjecture that this phenomenon arises from the inaccurate estimation of the precision structure; see Appendix A. To address this issue, we propose an algorithm that applies a more direct treatment to the precision structure compared with LNSMI.
@@ 3.2 Numerical algorithm of the proposed method
The quantities $V_{k}$ and $\Omega_{k}$ denote the estimated covariance and precision structures, respectively, computed at the $k$-th iteration. The specific procedure for determining the value of the penalty parameter is described in Subsection 5.2.
The proposed algorithm iterates the following three steps until a prescribed stopping criterion is satisfied. First, we estimate $\Omega$ by applying a shrinkage step of the same form as that used in the EM-type estimator, given $V_{k}$. Next, we update $V$ in an $M$-estimation manner, using the resulting $\Omega_{k+1}$. Then, we perform trace normalization so that $\mathrm{Tr}[V_{k+1}]=p$ holds. For the determination of the shrinkage coefficient and the stopping criterion, see Remark 8 and Subsection 5.2.
While LNSMI and RegTYL shrink the estimated covariance structure toward $I_{p}$ at each iteration, the proposed method shrinks the estimated precision structure toward $I_{p}$ at each iteration. In addition, the proposed method differs from HME in the way it shrinks the precision structure toward the identity matrix and in whether trace normalization is applied.
@@ 3.3 Interpretation of the proposed method as a penalized $M$-estimator
when $\det(I_{p}-\alpha\Omega^{-1})\neq 0$. Equation (8) represents the relation that the convergence point of the proposed algorithm formally satisfies for an appropriate choice of $\alpha$ and $w(\cdot)$. Furthermore, from the derivation of (8), it follows that $I_{p}-\alpha\Omega^{-1}\in\mathcal{S}_{++}^{p}$.
@@
Under the constraint $\mathrm{Tr}[\Omega]=c$ for a constant $c$, the term $\log\left|\det\left(I_{p}-\alpha V\right)\right|$ is minimized when $\Omega=(c/p)I_{p}$. This result is consistent with the shrinkage direction employed in the proposed algorithm.
@@
for $\alpha\in(0,\infty)$ is considered, the estimating equation is obtained as
$\Omega=\frac{1}{\alpha}\left\{\frac{1}{n}\sum_{i\in[n]}w(\bm{x}_{i}^{\top}\Omega\bm{x}_{i})\bm{x}_{i}\bm{x}_{i}^{\top}\right\}^{-1}+\frac{\alpha-1}{\alpha}I_{p}$
@@
3.4 Characterization of the proposed method by its estimating equation
@@
Alternatively, substituting $V^{-1}=(\Omega-\alpha I_{p})/(1-\alpha)$ into the estimating equation (8), leads to the more direct form
@@
When $\alpha=0$ (or $\alpha\searrow 0$) in the estimating equation (11), it reduces to the estimating equation (4), that is, the ordinary $M$-estimator. On the other hand, when $\alpha=1$, it corresponds to the class of GSSCMs [27]. In particular, when $w(s)=1/s$ ($s>0$), the SSCM is obtained.
@@ 3.5 Interpretation of the proposed method as a Bayesian estimator
We discuss the relationship between the proposed method and the maximum a posteriori estimation (MAP for short). Let $\bm{D}=(D_{1},D_{2},\ldots,D_{p})^{\top}$ be a random vector taking values in $\Delta^{p-1}$, and suppose that $\bm{D}\sim\textnormal{Dir}_{p}(\gamma_{1})$. The notation $\textnormal{Dir}_{p}(\gamma_{1})$ denotes the $p$-dimensional symmetric Dirichlet distribution with parameter $\gamma_{1}>0$. For $\gamma_{2}>0$, the random vector $\bm{\Lambda}=(\Lambda_{1},\Lambda_{2},\ldots,\Lambda_{p})^{\top}\in\mathbb{R}^{p}$ obtained through the variable transformation
@@
$f_{\bm{\Lambda}}(\bm{\lambda};\gamma_{1},\gamma_{2})=\frac{\Gamma\left(p\gamma_{1}\right)}{\left(\Gamma(\gamma_{1})\right)^{p}}\cdot\gamma_{2}^{p}\cdot\prod_{k=1}^{p}\left|1-\gamma_{2}\lambda_{k}\right|^{\gamma_{1}-1}\quad\left(\bm{\lambda}=(\lambda_{1},\ldots,\lambda_{p})^{\top}\in\Delta_{2}\right)\text{.}$
The support, $\Delta_{2}\coloneqq\{\bm{\eta}\in\mathbb{R}^{p};\,0\leq\eta_{k}\leq 1/\gamma_{2}\,(k\in[p]),\sum_{k}\eta_{k}=(p-1)/\gamma_{2}\}$.
Let $P_{\bm{\Lambda}}^{(\gamma_{1},\gamma_{2})}$ denote the distribution induced by the probability density function $f_{\bm{\Lambda}}(\bm{\cdot};\,\gamma_{1},\gamma_{2})$. We consider the hierarchical model with
@@
Focusing on the case $\gamma_{1}=n+1$ and, as in Subsection 2.4, allowing $\rho(\cdot)$ to be a general function not necessarily related to the density generator $g(\cdot)$, the problem of maximizing the log posterior reduces to the minimization of (7).
It can be seen that $\gamma_{2}>0$ in the prior distribution $P_{\bm{\Lambda}}^{(n+1,\gamma_{2})}$ ensures that the largest eigenvalue of the scatter matrix $V$ remains finite. It can be seen that $\gamma_{2}>0$ in the prior distribution $P_{\bm{\Lambda}}^{(n+1,\gamma_{2})}$ ensures that the largest eigenvalue of the scatter matrix $V$ remains finite.
3.6 Existence of the proposed method on $\mathcal{S}_{++}^{p}$
In this subsection, we discuss the existence of $V$ in $\mathcal{S}_{++}^{p}$ that satisfies the equations
@@ 3.6 Existence of the proposed method on $\mathcal{S}_{++}^{p}$
First, the existence in $\mathcal{S}_{++}^{p}$ of $V$ that satisfies (12) is established by our next theorem.
@@
From the proof of Theorem 3.1, we obtain a sequence $\{V_{k}\}_{k=0}^{\infty}$ of covariance structures, which is monotonically nonincreasing in the sense of the Löwner partial order. This is the reason why the proposed method incorporates a trace normalization step. Without trace normalization, the scale of the estimated covariance structure (for example, its trace) may approach zero, which could make the shrinkage factor practically meaningless. To avoid this issue, the proposed algorithm conducts trace normalization.
@@
It should be noted that Theorem 3.1, Corollary 3.1, and Lemma 3.1 can be discussed in the same manner for (13) and the estimating equation (11).
4 Robustness of proposed method
@@
Let Condition 2.1-(A)–(D) and (F) hold and suppose that $\kappa>n(p-1)/(n-p)$. If $\varepsilon_{m}<1-p/\kappa-(p-1)/n$, then $Z\in\mathcal{E}_{m}(X)$.
Let Condition 2.1-(B) and (C) hold and suppose that $\kappa>p$. If $\varepsilon_{m}>1-p/\kappa$, then $Z\notin\mathcal{E}_{m}(X)$.
The next lemma provides sufficient conditions on $\varepsilon_{m}$ under which $\lambda_{1}(\hat{V}(Z))$ and $1/\lambda_{p}(\hat{V}(Z))$ are bounded above, provided that $\hat{\theta}(\mathrm{Z})$ exists.
@@
Let Condition 2.1-(A)–(C) hold. If $\varepsilon_{m}=1$, then the set $\{\lambda_{1}(\hat{V}(Z));\,Z\in\mathcal{E}_{m}(X)\}$ is bounded above.
Let Condition 2.1-(B), (C), and (E) hold. If $\varepsilon_{m}<1-p/\kappa$, then the set $\{1/\lambda_{p}(\hat{V}(Z));\,Z\in\mathcal{E}_{m}(X)\}$ is bounded above.
@@
Lemma 4.2-(1) provides an important desirable property of the proposed estimator. The proposed estimator does not break down from above in the sense that its largest eigenvalue does not diverge to positive infinity when $\hat{V}(Z)$ exists.
@@
For non-penalized $M$-estimators, a stringent condition $\varepsilon_{m}<1/\kappa$ is sufficient to obtain the same conclusion [39] of Lemma 4.2-(1). In addition, for RegTYL, LNSMI, and HME, the condition $\kappa<1$ is sufficient for the same conclusion [40]. However, the condition $\kappa<1$ is incompatible with the necessary condition $\kappa\geq p$ for the existence of an non-penalized $M$-estimator in $\mathcal{S}_{++}^{p}$. Hence, RegTYL, LNSMI, and HME involve a trade-off between bias and robustness [40]. In contrast, the proof for the proposed method shows that the milder condition $\kappa>1$ is sufficient.
@@
$1-\frac{p}{\kappa}-\frac{p-1}{n}\leq\delta^{*}(X;\,\hat{V})\leq 1-\frac{p}{\kappa}.$
@@
The upper bound $1-p/\kappa$ of the breakdown point in Theorem 4.1 is attained when there exists a point mass at the origin.

### v2

@@ 2.4 Penalized $M$-estimation
which is the negative log-likelihood divided by $n$. The function $\rho^{*}(s)\coloneqq-\log g(s)$ is defined for $s\geq 0$.
@@
and seek $V\in\mathcal{S}_{++}^{p}$ that minimizes the objective function (2). In this case, the minimizer $V$ of (2) is an $M$-estimator [19, 22, 35]. For example, TME corresponds to the choice $\rho(s)=p\log s$ ($s>0$) in (2) [35]. In this paper, no assumptions are imposed on the function $\rho(\cdot)$ itself. Instead, assumptions are placed on the function corresponding to its first derivative, as stated in Condition 2.1.
@@ 2.5 Characterization of the estimator
A solution satisfying the estimating equation (4) is linearly equivariant (see Remark 2), whereas a solution satisfying (5) is not, and this difference affects the robustness of the estimator [6]. In (3), when the Kullback–Leibler type penalty
@@
with $\alpha\in(0,\infty)$ is considered, the estimating equation is obtained as
@@
The $M$-estimator of $V$ is often defined through estimating equations such as (4), (5), and (6), and computed using methods based on fixed-point algorithms for the corresponding estimating equations. Estimators characterized as solutions to the estimating equations (4), (5), and (6) do not necessarily possess scale equivariance. See Remark 2 for the definition of scale equivariance.
@@
    The weight function $w:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ is continuous and nonincreasing.
@@
    For $\psi(s)\coloneqq sw(s)$ ($s\geq 0$), it holds that $\kappa\coloneqq\sup_{s\geq 0}\{\psi(s)\}<\infty$.
@@
    The function $\psi:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ is strictly increasing in $\{s\in\mathbb{R}_{\geq 0}:\psi(s)<\kappa\}$.
@@
    For any (p1)(p-1)-dimensional linear subspace $H\subset\mathbb{R}^{p}$, it holds that $P(H)<1-p/\kappa$.
@@
for $c>0$, the $t$-distribution type weight functions $w(s)=(p+\nu)/(s+\nu)$ ($s\geq 0$) for $\nu\in\mathbb{N}$, and the Tyler type weight function $w(s)=p/s$ ($s\geq 0$).
2.6 Additive finite sample breakdown point
The breakdown point is one of the measures of an estimator’s robustness against contamination by outliers in the data [13, 19, 20, 38, 39]. Several definitions of the breakdown point exist, which can be broadly classified into those in the asymptotic setting and those in the finite sample setting. In this paper, we focus on the breakdown point in the finite sample setting.
@@ 2.6 Additive finite sample breakdown point
where $\mathcal{E}_{m}(\mathrm{X})=\{\mathrm{Z}=\mathrm{X}\cup\mathrm{Y}:\,\hat{\theta}(\mathrm{Z})\textnormal{ exists in }\Theta\,(\mathrm{X}:\text{fixed})\}$.
@@
$\varepsilon^{*}\left(\hat{V}(\cdot),\mathrm{X}\right)\coloneqq\min\left\{\varepsilon_{m};\,b\left(\varepsilon_{m};\,\hat{V}(\cdot),\mathrm{X}\right)=\infty\right\}.$
We consider a discrepancy function $B(V_{1},V_{2})$ that satisfies
$B\left(V_{1},V_{2}\right)\to\infty\quad\text{as }\lambda_{p}(V_{2})\to 0\ \text{or}\ \lambda_{1}(V_{2})\to\infty.$
Examples of such discrepancy functions include $\mathrm{Tr}[V_{1}V_{2}^{-1}+V_{1}^{-1}V_{2}]$ and $\|\log(V_{1}^{-1/2}V_{2}V_{1}^{-1/2})\|_{\mathrm{F}}$ [38, 39]. Here, for $X\in\mathcal{S}_{++}^{p}$, we define
@@
where $Q\,\mathrm{diag}\left(\lambda_{1}(X),\dots,\lambda_{p}(X)\right)\,Q^{\top}$ is the eigenvalue decomposition of $X$ with an orthogonal matrix $Q$.
@@ 3.1 Limitations of existing methods
We examine the performance of TME and LNSMI through numerical experiments under two types of outlier configurations. We observe that when outliers form clusters, the performance of TME and LNSMI deteriorates significantly for a sample size of $N\coloneqq n+m=100$ and an outlier ratio of $3\%$, when the data dimension exceeds $p\approx 35$. See Section 5 for the detailed settings and Figure 3-(b, d, f) for the results. Moreover, this phenomenon is found to be closely related to the breakdown point of TME. Based on the results of the numerical experiments, we conjecture that this phenomenon arises from the inaccurate estimation of the precision structure; see Appendix A. To address this issue, we propose an algorithm that applies a more direct treatment of the precision structure compared with LNSMI.
@@ 3.2 Numerical algorithm of the proposed method
The quantities $V_{k}$ and $\Omega_{k}$ denote the estimated covariance structure and precision structure, respectively, computed at the $k$-th iteration. The specific procedure for determining the value of the penalty parameter is described in Subsection 5.2.
The proposed algorithm iterates the following three steps until a prescribed stopping criterion is satisfied. First, we estimate $\Omega$ by applying a shrinkage step of the same form as that used in the EM-type estimator, given $V_{k}$. Next, we update $V_{k+1}$ in an $M$-estimation manner, using the resulting $\Omega_{k+1}$. Then, we perform trace normalization so that $\mathrm{Tr}[V_{k+1}]=p$ holds. For the determination of the shrinkage coefficient and the stopping criterion, see Remark 8 and Subsection 5.2.
While LNSMI and RegTYL shrink the estimated covariance structure toward $I_{p}$ at each iteration, the proposed method shrinks the estimated precision structure toward $I_{p}$ at each iteration. In addition, the proposed method differs from HME in the way it shrinks the precision structure toward $I_{p}$ and in whether trace normalization is applied.
@@ 3.3 Interpretation of the proposed method as a penalized $M$-estimator
when $\det(I_{p}-\alpha\Omega^{-1})\neq 0$. Equation (8) represents the relation formally satisfied by the convergence point of the proposed algorithm for an appropriate choice of $\alpha$ and $w(\cdot)$. Furthermore, from the derivation of (8), it follows that $I_{p}-\alpha\Omega^{-1}\in\mathcal{S}_{++}^{p}$.
@@
Under the constraint $\mathrm{Tr}[V]=c$ for a constant $c$, the term $\log\left|\det\left(I_{p}-\alpha V\right)\right|$ is minimized when $V=(c/p)I_{p}$. This result is consistent with the shrinkage direction employed in the proposed algorithm.
@@
with $\alpha\in(0,\infty)$ is considered, the estimating equation is obtained as
$\Omega=\frac{1}{\alpha}\left\{\frac{1}{n}\sum_{i=1}^{n}w(\bm{x}_{i}^{\top}\Omega\bm{x}_{i})\bm{x}_{i}\bm{x}_{i}^{\top}\right\}^{-1}+\frac{\alpha-1}{\alpha}I_{p}$
@@
3.4 Characterization of the proposed method by its estimating equations
@@
Alternatively, substituting $V^{-1}=(\Omega-\alpha I_{p})/(1-\alpha)$ into the estimating equation (8) leads to the more direct form
@@
When $\alpha\searrow 0$, the estimating equation (11) coincides with (4), namely, the ordinary $M$-estimator. On the other hand, when $\alpha=1$, the resulting estimator corresponds to the class of GSSCMs [26]. In particular, when $\alpha=1$ and $w(s)=1/s$ ($s>0$), the SSCM is obtained as a special case.
@@ 3.5 Interpretation of the proposed method as a Bayesian estimator
We discuss the relationship between the proposed method and the maximum a posteriori estimation (MAP for short). Let $\bm{D}=(D_{1},D_{2},\ldots,D_{p})^{\top}$ be a random vector taking values in $\Delta^{p-1}$, and suppose that $\bm{D}\sim\textnormal{Dir}_{p}(\gamma_{1})$. The notation $\textnormal{Dir}_{p}(\gamma_{1})$ denotes the $p$-dimensional symmetric Dirichlet distribution with parameter $\gamma_{1}>0$. For $\gamma_{2}>0$, the random vector $\bm{\Lambda}=(\Lambda_{1},\Lambda_{2},\ldots,\Lambda_{p})^{\top}\in\mathbb{R}^{p}$ obtained through the linear transformation
@@
$f_{\bm{\Lambda}}(\bm{\lambda};\gamma_{1},\gamma_{2})=\frac{\Gamma\left(p\gamma_{1}\right)}{\left(\Gamma(\gamma_{1})\right)^{p}}\cdot\gamma_{2}^{p}\cdot\prod_{k=1}^{p}\left|1-\gamma_{2}\lambda_{k}\right|^{\gamma_{1}-1}\quad\left(\bm{\lambda}=(\lambda_{1},\ldots,\lambda_{p})^{\top}\in\Delta_{2}\right)\text{,}$
where the support is $\Delta_{2}\coloneqq\{\bm{\eta}\in\mathbb{R}^{p}:0\leq\eta_{k}\leq 1/\gamma_{2}\,(k\in[p]),\sum_{k}\eta_{k}=(p-1)/\gamma_{2}\}$.
Let $P_{\bm{\Lambda}}^{(\gamma_{1},\gamma_{2})}$ denote the distribution induced by the probability density function $f_{\bm{\Lambda}}(\cdot;\,\gamma_{1},\gamma_{2})$. We consider the hierarchical model given by
@@
Focusing on the case $\gamma_{1}=n+1$ and, as in Subsection 2.4, allowing $\rho(\cdot)$ to be a general function not necessarily related to the density generator $g(\cdot)$, the problem of maximizing the log posterior reduces to the minimization of the objective function (7).
The condition $\gamma_{2}>0$ in the prior distribution $P_{\bm{\Lambda}}^{(n+1,\gamma_{2})}$ ensures that the largest eigenvalue of the scatter matrix $V$ remains finite.
3.6 Existence of the proposed estimator in $\mathcal{S}_{++}^{p}$
In this subsection, we discuss the existence of $V\in\mathcal{S}_{++}^{p}$ that satisfies the following equations
@@ 3.6 Existence of the proposed estimator in $\mathcal{S}_{++}^{p}$
First, our next theorem establishes the existence of a solution $V\in\mathcal{S}_{++}^{p}$ to (12).
@@
The proof of Theorem 3.1 shows that the sequence of covariance structures obtained from the iteration without the trace normalization step is monotonically nonincreasing with respect to the Löwner partial order. This explains why the proposed algorithm incorporates a trace normalization step. Without trace normalization, the scale of the estimated covariance structure (for example, its trace) may approach zero, which could make the shrinkage factor practically meaningless. To avoid this issue, the proposed algorithm conducts trace normalization.
@@
Theorem 3.1, Corollary 3.1, and Lemma 3.1 can be derived in the same manner for (13) and (11).
4 Robustness of the proposed estimator
@@
Let Condition 2.1-(A)–(D) and (F) hold and suppose that $\kappa>n(p-1)/(n-p)$. If $\varepsilon_{m}<1-p/\kappa-(p-1)/n$, then $\mathrm{Z}\in\mathcal{E}_{m}(\mathrm{X})$.
Let Condition 2.1-(B) and (C) hold and suppose that $\kappa>p$. If $\varepsilon_{m}>1-p/\kappa$, then $\mathrm{Z}\notin\mathcal{E}_{m}(\mathrm{X})$.
The next lemma provides sufficient conditions on $\varepsilon_{m}$ under which $\lambda_{1}(\hat{V}(\mathrm{Z}))$ and $1/\lambda_{p}(\hat{V}(\mathrm{Z}))$ are bounded above, provided that $\hat{V}(\mathrm{Z})$ exists.
@@
Let Condition 2.1-(A)–(C) hold. If $\varepsilon_{m}=1$, then the set $\{\lambda_{1}(\hat{V}(\mathrm{Z}));\,\mathrm{Z}\in\mathcal{E}_{m}(\mathrm{X})\}$ is bounded above.
Let Condition 2.1-(B), (C), and (E) hold. If $\varepsilon_{m}<1-p/\kappa$, then the set $\{1/\lambda_{p}(\hat{V}(\mathrm{Z}));\,\mathrm{Z}\in\mathcal{E}_{m}(\mathrm{X})\}$ is bounded above.
@@
Lemma 4.2-(1) provides a desirable property of the proposed estimator. The proposed estimator does not breakdown from above, in the sense that its largest eigenvalue remains bounded whenever $\hat{V}(\mathrm{Z})$ exists.
@@
For non-penalized $M$-estimators, a stringent condition $\varepsilon_{m}<1/\kappa$ is sufficient to obtain the same conclusion as in Lemma 4.2-(1) [38]. In addition, for RegTYL, LNSMI, and HME, the condition $\kappa<1$ is sufficient for the same conclusion [39]. However, the condition $\kappa<1$ is incompatible with the necessary condition $\kappa\geq p$ for the existence of a non-penalized $M$-estimator in $\mathcal{S}_{++}^{p}$. Hence, RegTYL, LNSMI, and HME involve a trade-off between bias and robustness [39]. In contrast, the proof for the proposed estimator shows that the much milder condition $\kappa>1$ is sufficient.
@@
$1-\frac{p}{\kappa}-\frac{p-1}{n}\leq\varepsilon^{*}\left(\hat{V}(\cdot),\mathrm{X}\right)\leq 1-\frac{p}{\kappa}.$
@@
The upper bound $1-p/\kappa$ of the AdBP in Theorem 4.1 is attained when there exists a point mass at the origin.